I need help with this equation

n-(8+4n)=2(3n-4)

Answer by gonzo(654)

equation you want to solve is
n-(8+4n) = 2(3n-4)
this translates to
n - (8+4*n) = 2*(3*n-4)
the equation then becomes
n - 8 - 4*n = 2*3*n - 2*4
this becomes
n - 8 - 4*n = 6*n - 8
adding 8 to both sides of the equation, we get
n - 4*n = 6*n
this becomes
-3*n = 6*n
adding 3*n to both sides of the equation, we get
0 = 6*n + 3*n
this becomes
0 = 9*n
dividing both sides of the equation by 9, we get
0 = n
this is the same as
n = 0